# The Cone Restriction: An old Approach Revisited

### v5

@@ 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
Cellular-dominant case Suppose
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.6)
Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Now
$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right),\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right)+1,\ \ A_{j+1}=A_{j}$
$C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{\delta}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
$C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
Also, let ${\rho}_{j+1}:={\rho}_{j}/2$. Since the cellular case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}$, there exists a polynomial $P$ with degree $O\left(d\right)$ such that
1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
2. For all $\in\mathcal{O}, OO$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
3. ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O\right)}\sim d^{-m}{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)},\ \ \forall O\in\mathcal{O}$
where $\mathcal{O}\left(O_{j}\right)$ is a refinement of the set of connected components of $O_{j}\backslash N_{{\rho}^{1/2+{\delta}_{m}}_{j}}Z\left(P\right)$. For $O\in\mathcal{O}\left(O_{j}\right)$, define
$f_{O}:=\sum_{\begin{array}[]{c}\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]\\ T^{l}_{\theta,v}\cap O\neq\emptyset\end{array}}{{\left(f_{O_{j}}\right)}^{l}_{\theta,v}}$
By Lemma 2.30, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus Lemma 2.1 gives
$\sum_{O\in\mathcal{O}\left(O_{j}\right)}{{\left\|f_{O}\right\|}^{2}_{2}}\lesssim d{\left\|f_{O_{j}}\right\|}^{2}_{2}$
By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with ${\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
${\left\|f_{O}\right\|}^{2}_{2}\lesssim d^{-m+1}{\left\|f_{O_{j}}\right\|}^{2}_{2},\ \ \forall O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$
Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\mathcal{B}\ \sim\ 1$ such that
$N_{{\rho}^{1/2+{\delta}_{m}}_{j}}{\boldsymbol{\mathrm{Z}}}_{O_{j}}\subseteq\bigcup_{b\in\mathcal{B}}{N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right)}$
For $O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$ and $b\in\mathcal{B}$, define
$O_{b}:=O\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right),\ \ \widetilde{f_{O_{b}}}:=\sum_{\begin{array}[]{c}\left(\zeta,w,L\right)\in\mathbb{T}\left[{\rho}_{j+1}\right]\\ T^{L}_{\zeta,w}\left(y_{O}\right)\cap O_{b}\neq\emptyset\end{array}}{{\left(\widetilde{f_{O}}\right)}^{L}_{\zeta,w}}$
here $\widetilde{f_{O}}\left(\xi\right):=e^{i{\psi}_{y_{O}}\left(\xi\right)}f_{O}\left(\xi\right)$. By Lemma 2.6, $f_{O_{b}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b,B_{O}\right)$. Let
${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{\bigcup_{O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)}{\left\{O_{b}:\ b\in\mathcal{B}\right\}}}$
Now we verify the properties. For Property 1, 2, and 3, see [11] for details. Here we verify Property 4. By Lemma 2.1 and 2.6, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$,
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.7)
where $\uparrow\mathcal{T}$ denotes set the $\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]$ such that
@@ 1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-\frac{1}{2}}_{j+1},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j+1}}\right),\ T^{l}_{\theta,v}\cap B\left(y_{O_{j+1}},{\rho}_{j+1}\right)\right)\leq C{\rho}^{1/2+\delta}_{j}$
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in Property 4. Combine (3.5) and (3.7),
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
Since ${\rho}_{j+1}\ \sim\ {\rho}_{j}$, by choosing $d$ large enough, one has
@@
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
[…]
@@
Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Now
@@
Let ${\rho}_{j+1}:={\rho}^{1-{\widetilde{\delta}}_{m-1}}_{j}$. Since the algebraic case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, there exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq{\boldsymbol{\mathrm{Z}}}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\right)}$
where $C_{\mathrm{alg}}$ is large constant. Cover $O_{j}$ by finitely overlapping ${\rho}_{j+1}$-balls $B$. Let ${\mathcal{B}}_{O_{j}}$ denotes the set of such balls. For each $B\in{\mathcal{B}}_{O_{j}}$, let
${\mathbb{T}}_{B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]:T^{l}_{\theta,v}\cap B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\neq\emptyset\right\},f_{B}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B}}$
Let ${\mathbb{T}}_{B\mathrm{,tang}}\subseteq{\mathbb{T}}_{B}$ be the set of ${\rho}_{j}$-scale wave packets $\left(\theta,v,l\right)$ such that
1. $T^{l}_{\theta,v}\cap 2B\subseteq N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\boldsymbol{=}N_{{\rho}^{1/2+{\delta}_{m-1}}_{j+1}}\boldsymbol{\mathrm{Y}}$
2. If $x\in T^{l}_{\theta,v}$ and $y\in\boldsymbol{\mathrm{Y}}\mathrm{\cap}2B$ satisfy $\left|y-x\right|\lesssim{\rho}^{1/2+{\delta}_{m}}_{j}=\ {\rho}^{1/2+{\delta}_{m-1}}_{j+1}$, then
$\angle\left(L\left(\theta\right),T_{y}\boldsymbol{\mathrm{Y}}\right)\lesssim{\rho}^{-1/2+{\delta}_{m-1}}_{j-1}$
Let ${\mathbb{T}}_{B\mathrm{,trans}}:={\mathbb{T}}_{B}\backslash{\mathbb{T}}_{B\mathrm{,tang}}$ and
$f_{B,\mathrm{tang}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,tang}}},\ \ f_{B,\mathrm{trans}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,trans}}}$
@@
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}$
$\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
By pigeonholing, at least one of the following situations holds.
Tangential subcase: Let $C_{\mathrm{tang}}$ be a large constant. If
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq\left(C_{\mathrm{tang}}/2\right)\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$ (3.9)
then the algorithm terminates, and we set $J=j+1$. We mark this situation as [tang]. Let $\mathcal{S}$ be a collection of grains, which is defined by
$\mathcal{S}:=\left\{\left(\boldsymbol{\mathrm{Y}},B\right):B\in{\mathcal{B}}_{O_{j}}\mathrm{\ for\ some\ }O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}\right\}$
For $S=\left(\boldsymbol{\mathrm{Y}},B\right)\in\mathcal{S}$, define $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Then Lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, (3.9) gives
$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B_{S}\right)}}$ (3.10)
By Lemma 2.1 and Corollary 2.3, for all $1\leq\rho\leq{\rho}_{j+1}$, there holds

### v7

@@ 2.8. Polynomial Partitioning
Finally, we set
In [12, 8, 10], a result on nested tubes as a corollary of Lemma 2.11 in [25]—was used to obtain new results on the Bochner–Riesz problem, Stein’s square function, and Fourier restriction for the paraboloid. We state this result below.
Lemma 2.26 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
Consequently,
Corollary 2.27.
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
Proof.
3. Finding Polynomial Structure
In this section, we reformulate the induction arguments of [15] as a recursive procedure, combining ideas from [11, 12]. The argument consists of two algorithms.
In Algorithm 1, we iteratively apply polynomial partitioning until the tangential term dominates. This procedure reduces the problem from an $m$-dimensional to an (m1)(m-1)-dimensional setting and corresponds to the induction on scale. Algorithm 2 is obtained by repeatedly applying Algorithm 1, and the iteration terminates once the minimal dimension is reached.
Throughout the iteration, we assume without loss of generality that the error terms $\mathrm{RapDec}(R)\|f\|_{2}$ are negligible.
3.1. Algorithm 1
This algorithm proceeds via iterated polynomial partitioning. At each step, one of three cases occurs: the cellular case, the transverse case, or the tangential case. The iteration terminates either when the tangential case occurs or when the scale becomes sufficiently small.
Input. Algorithm 1 takes the following data as input:
  • A grain $(\mathbf{Z},B(y,r))$ of dimension $m$.
  • A function $f\in L^{\infty}\!\left(2\overline{B}^{\,n-1}\setminus B^{n-1}\right)$ that is tangent to $(\mathbf{Z},B(y,r))$.
  • An admissible integer $A\in\mathbb{N}$.
At the initial step, we set
$O_{0}:=\{\,B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\,\},\qquad f_{O_{0}}:=f,\qquad\mathfrak{h}_{0}:=\emptyset,\qquad\rho_{0}:=r.$
Output. At the $j$-th iteration, Algorithm 1 outputs:
  • History. A word $\mathfrak{h}_{j}$ of length $j$ in the alphabet $\{\mathrm{a},\mathrm{c}\}$, recording how each cell $O_{j}\in\mathcal{O}_{j}$ is produced by the iterated polynomial partitioning. Let $\mathrm{a}(j)$ and $\mathrm{c}(j)$ denote the number of letters $\mathrm{a}$ and $\mathrm{c}$ in $\mathfrak{h}_{j}$, respectively.
  • Spatial scale $\rho_{j}$. The scale $\rho_{j}$ is determined by the initial scale $r$ and the history $\mathfrak{h}_{j}$. Define an auxiliary parameter $\widetilde{\delta}_{m-1}$ by
    $(1-\widetilde{\delta}_{m-1})\!\left(\frac{1}{2}+\delta_{m-1}\right)=\frac{1}{2}+\delta_{m},$
    which implies
    $\frac{\delta_{m-1}}{2}\leq\widetilde{\delta}_{m-1}\leq 2\delta_{m-1}.$
  • Cells. A family of subsets $\mathcal{O}_{j}\subset\mathbb{R}^{n}$. Each cell $O_{j}\in\mathcal{O}_{j}$ is contained in a ball $B_{O_{j}}=B(y_{O_{j}},\rho_{j})$.
  • Localized functions. A collection of functions $\{f_{O_{j}}\}_{O_{j}\in\mathcal{O}_{j}}$. For each cell $O_{j}$ there exists a translate $\mathbf{Z}_{O_{j}}:=\mathbf{Z}+x_{O_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $(\mathbf{Z}_{O_{j}},B_{O_{j}})$.
  • Auxiliary parameters. A large integer $d$ depending only on the admissible parameters and $\deg\mathbf{Z}$, together with
    $C^{\mathrm{I}}_{j,\delta}(d,r):=d^{\#\mathrm{c}(j)\delta}\,(\log r)^{\#\mathrm{a}(j)(1+\delta)},\qquad C^{\mathrm{II}}_{j,\delta}(d):=d^{\#\mathrm{c}(j)\delta+n\#\mathrm{a}(j)(1+\delta)},$
    $C^{\mathrm{III}}_{j,\delta}(d,r):=d^{j\delta}r^{\overline{C}\#\mathrm{a}(j)\delta_{m}},\qquad C^{\mathrm{IV}}_{j,\delta}(d,r):=C^{\mathrm{III}}_{j,\delta}(d,r),\qquad A_{j}:=3^{-\#\mathrm{a}(j)}A.$
Properties. At the $j$-th iteration:
  1. (1)
    The mass of $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}$ is concentrated on the cells $O_{j}\in\mathcal{O}_{j}$:
    $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}\leq C^{\mathrm{I}}_{j,\delta}(d,r)\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.1)
  2. (2)
    The functions $f_{O_{j}}$ satisfy
    $\sum_{O_{j}\in\mathcal{O}_{j}}\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{II}}_{j,\delta}(d)\,d^{\#\mathrm{c}(j)}\|f\|_{2}^{2}.$ (3.2)
  3. (3)
    Each $f_{O_{j}}$ satisfies, for all $1\leq\rho\leq\rho_{j}$,
    $\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j)(m-1)}\|f\|_{2}^{2},$ (3.3)
    and for all $\rho\in[1,\rho_{j}]$,
@@ 3.1. Algorithm 1
    $\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j}}\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}\leq\\ C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}$ (3.4)
  4. (4)
    For $\mathcal{T}\subseteq\mathbb{T}[\rho_{j}]$ and $O_{j}\in\mathcal{O}_{j}$,
    $\|f_{O_{j},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}\mathcal{T}}\|_{2}^{2},$ (3.5)
    where $(f_{O_{j},\mathcal{T}})^{\widetilde{}}:=({f_{O_{j}}})^{\widetilde{}}_{\mathcal{T}}$, and $\uparrow^{j}\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[r]$ such that
    $\mathrm{dist}_{H}(\theta,\zeta)\leq c_{j}\rho_{j}^{-1/2}$
    $\mathrm{dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j}}),\,T_{\theta,v}^{l}(y)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
    for some $(\zeta,w,L)\in\mathcal{T}$, where
    $c_{j}:=\sum_{i=1}^{j}2^{-(j-1)/2}C$
    for an absolute constant $C$.
Recursion Steps. If $\rho_{j}\leq r^{\widetilde{\delta}_{m-1}}$, then the algorithm terminates. In this case we set $J:=j$ and $\mathcal{O}:=\mathcal{O}_{j}$, and label this outcome as [tiny]. Otherwise, for each $O_{j}\in\mathcal{O}_{j}$ we apply polynomial partitioning with degree $d$ to $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}$.
Let $\mathcal{O}_{j,\mathrm{cell}}\subseteq\mathcal{O}_{j}$ denote the collection of cells for which the cellular case occurs, and set
$\mathcal{O}_{j,\mathrm{alg}}:=\mathcal{O}_{j}\setminus\mathcal{O}_{j,\mathrm{cell}}.$
then one of the following cases holds:
Cellular-dominant case. Suppose
$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}\leq\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.6)
Append the letter $\mathrm{c}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Then
$\mathrm{a}(j+1)=\#\mathrm{a}(j),\qquad\#\mathrm{c}(j+1)=\#\mathrm{c}(j)+1,\qquad A_{j+1}=A_{j},$
and
$C^{\mathrm{I}}_{j+1,\delta}=d^{\delta}C^{\mathrm{I}}_{j,\delta},\qquad C^{\mathrm{II}}_{j+1,\delta}=d^{\delta}C^{\mathrm{II}}_{j,\delta},$
$C^{\mathrm{III}}_{j+1,\delta}=d^{\delta}C^{\mathrm{III}}_{j,\delta},\qquad C^{\mathrm{IV}}_{j+1,\delta}=d^{\delta}C^{\mathrm{IV}}_{j,\delta}.$
Set $\rho_{j+1}:=\rho_{j}/2$. For each $O_{j}\in\mathcal{O}_{j,\mathrm{cell}}$, the cellular case yields a polynomial $P$ of degree $O(d)$ such that:
  1. (1)
    $\mathcal{O}(O_{j})\sim d^{m}$,
  2. (2)
    every $O\in\mathcal{O}(O_{j})$ lies in a ball $B_{O}=B(y_{O},\rho_{j+1})$,
  3. (3)
    $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O)}\sim d^{-m}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}$.
Here $\mathcal{O}(O_{j})$ refines the connected components of $O_{j}\setminus N_{\rho_{j}^{1/2+\delta_{m}}}Z(P)$. Define
$f_{O}:=\sum_{\begin{subarray}{c}(\theta,v,l)\in\mathbb{T}[\rho_{j}]\\ T_{\theta,v}^{l}\cap O\neq\emptyset\end{subarray}}(f_{O_{j}})_{\theta,v}^{l}.$
By Lemma 2.21 and Lemma 2.1,
$\sum_{O\in\mathcal{O}(O_{j})}\|f_{O}\|_{2}^{2}\lesssim d\,\|f_{O_{j}}\|_{2}^{2}.$
Pigeonholing gives a refinement $\mathcal{O}^{\prime}(O_{j})$ with $\mathcal{O}^{\prime}\sim d^{m}$ such that
$\|f_{O}\|_{2}^{2}\lesssim d^{-m+1}\|f_{O_{j}}\|_{2}^{2}.$
Since $\rho_{j+1}\sim\rho_{j}$, there exists a finite set $\mathcal{B}\subset B(0,\rho_{j}^{1/2+\delta_{m}})$ with $\mathcal{B}\sim 1$ such that
$N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Z}_{O_{j}}\subset\bigcup_{b\in\mathcal{B}}N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b).$
For $O\in\mathcal{O}^{\prime}(O_{j})$ and $b\in\mathcal{B}$ define
$O_{b}:=O\cap N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b),$
$\widetilde{f_{O_{b}}}:=\sum_{\begin{subarray}{c}(\zeta,w,L)\in\mathbb{T}[\rho_{j+1}]\\ T_{\zeta,w}^{L}(y_{O})\cap O_{b}\neq\emptyset\end{subarray}}(\widetilde{f_{O}})_{\zeta,w}^{L},\qquad\widetilde{f_{O}}(\xi)=e^{i\psi_{y_{O}}(\xi)}f_{O}(\xi).$
By Lemma 2.6, $f_{O_{b}}$ is tangent to $(\mathbf{Z}_{O_{j}}+b,B_{O})$. Set
$\mathcal{O}_{j+1}:=\bigcup_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\bigcup_{O\in\mathcal{O}^{\prime}(O_{j})}\{O_{b}:\,b\in\mathcal{B}\}.$
Properties 1–3 follow from [11]. We verify Property 4. By Lemma 2.1 and Lemma 2.6, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$,
$\|\widetilde{f}_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim\|f_{O_{j},\uparrow\mathcal{T}}\|_{2}^{2}.$ (3.7)
Here $\uparrow\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[\rho_{j}]$ such that
@@
$\mathrm{Dist}_{H}(\theta,\zeta)\leq C\rho_{j+1}^{-1/2},\qquad\mathrm{Dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j+1}}),\,T_{\theta,v}^{l}\cap B(y_{O_{j+1}},\rho_{j+1})\right)\leq C\rho_{j}^{1/2+\delta}$
for some $(\zeta,w,L)\in\mathcal{T}$, where $C$ is the absolute constant from Property 4.Combining (3.5) and (3.7), we obtain
$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
Since $\rho_{j+1}\sim\rho_{j}$, choosing $d$ sufficiently large yields
$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j+1}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
From [12] we see ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$. This completes the verification.
Algebraic-dominant Case. If (3.6) fails, then
@@
Append the letter $\mathrm{a}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Now we have
@@
Set $\rho_{j+1}:=\rho_{j}^{\,1-\widetilde{\delta}_{m-1}}$. Since the algebraic case occurs, for each $O_{j}\in\mathcal{O}_{j,\mathrm{alg}}$ there exists an (m1)(m-1)-dimensional transverse complete intersection $\mathbf{Y}\subseteq\mathbf{Z}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\bm{\mathrm{Y}}\right)}$
where $C_{\mathrm{alg}}$ is a sufficiently large constant. Let $\mathcal{B}_{O_{j}}$ denote the collection of such balls. For each $B\in\mathcal{B}_{O_{j}}$, define
$\mathbb{T}_{B}:=\left\{(\theta,v,l)\in\mathbb{T}[\rho_{j}]:T_{\theta,v}^{l}\cap B\cap N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}\neq\emptyset\right\},\qquad f_{B}:=(f_{O_{j}})_{\mathbb{T}_{B}}.$
Let $\mathbb{T}_{B,\mathrm{tang}}\subseteq\mathbb{T}_{B}$ be the set of $\rho_{j}$-scale wave packets $(\theta,v,l)$ satisfying:
  1. (1)
    $T_{\theta,v}^{l}\cap 2B\subseteq N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}=N_{\rho_{j+1}^{1/2+\delta_{m-1}}}\mathbf{Y}$;
  2. (2)
    whenever $x\in T_{\theta,v}^{l}$ and $y\in\mathbf{Y}\cap 2B$ satisfy $|x-y|\lesssim\rho_{j}^{1/2+\delta_{m}}=\rho_{j+1}^{1/2+\delta_{m-1}}$, then
    $\angle\!\big(L(\theta),T_{y}\mathbf{Y}\big)\lesssim\rho_{j+1}^{-1/2+\delta_{m-1}}.$
Define $\mathbb{T}_{B,\mathrm{trans}}:=\mathbb{T}_{B}\setminus\mathbb{T}_{B,\mathrm{tang}}$ and
$f_{B,\mathrm{tang}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{tang}}},\qquad f_{B,\mathrm{trans}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{trans}}}.$
@@
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
By a pigeonholing argument, at least one of the following cases holds.
Tangential Subcase. Let $C_{\mathrm{tang}}$ be a large constant. If
$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq\frac{C_{\mathrm{tang}}}{2}\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\sum_{B\in\mathcal{B}_{O_{j}}}\|Ef_{B,\mathrm{tang}}\|^{p}_{BL^{p}_{k,A_{j+1}}(B)},$ (3.9)
then the algorithm terminates. We set $J=j+1$ and label this outcome [tang]. Define the collection of grains
$\mathcal{S}:=\left\{(\mathbf{Y},B):B\in\mathcal{B}_{O_{j}}\ \text{for some}\ O_{j}\in\mathcal{O}_{j,\mathrm{alg}}\right\}.$
For each $S=(\mathbf{Y},B)\in\mathcal{S}$, set $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, inequality (3.9) yields
$\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}\|Ef_{S}\|^{p}_{BL^{p}_{k,A_{j+1}}(B_{S})}.$ (3.10)
By Lemma 2.1 and Corollary 2.3, for all $1\leq\rho\leq\rho_{j+1}$, we have